# CardioComposer: Leveraging Differentiable Geometry for Compositional Control of Anatomical Diffusion Models

### v2

@@ 5.3 Evaluation Metrics
We evaluate pairwise conditional fidelity for size, shape, and position by taking the $L_{1}$-norm between the target and sample moments. We measure morphological quality metrics by comparing the distribution of real and synthetic anatomy in morphological feature space kadry2024probing. To embed each label map, we consider all 10 tissues as substructures and concatenate, over all substructures, the masses, centroids, and eigenvalues of the normalized covariance matrices. We specifically use morphological variants of improved precision and recall, as well as the Fréchet distance (FD) kynkaanniemi2019improved; kadry2024probing. Lastly, we leverage pointcloud-based metrics to assess 3D shape yang2019pointflow, such as minimum matching distance (MMD), coverage (COV), and 1-nearest neighbor accuracy (1-NNA). Distances between pointclouds are computed with Earth Mover’s Distance (EMD). Further details can be found in Section 8.6.
@@ 5.4 Evaluating Anatomical Generation Quality
We first aim to compare and evaluate geometric control methods on both conditional fidelity and synthetic anatomy quality. We sample target moments for a single substructure (myocardium) from the validation set and generate 200 anatomical segmentations per method. We sweep over guidance weights $w\in[0,2]$. In Figure 4, we show that geometric guidance enhances conditional fidelity, especially at higher guidance weights. We observe that our method maintains generation quality, retaining similar levels of morphological and pointcloud evaluation metrics with increasing guidance. Figure 3 shows example label maps generated through varying guidance values for all methods, across guidance values for all methods; only our method and implicit conditioning align the target and sample ellipsoids under guidance. Further information on which features were plotted can be found in Section 8.13.
@@ 5.5 Evaluating Geometric Disentanglement
We next show that our guidance framework uniquely enables disentangled control of geometric attributes. We use 100 target moments for myocardial labels from the validation set using no losses (Uncond.), a combination of all losses ($\mathcal{L}_{\text{geom}}$), or each individual moment loss ($\mathcal{L}_{\text{size}},\mathcal{L}_{\text{pos}}$, and $\mathcal{L}_{\text{shape}}$). Figure 9 shows that each individual loss improves its corresponding conditional-fidelity metric while leaving the others approximately unchanged. The main exception is the interaction between shape and mass, where adding a guidance weight for the shape loss enhances mass fidelity. This phenomenon is likely due to correlation of size and shape in the dataset. Qualitative results can be seen in Figures 5 and 6, where the right ventricle is constrained independently by mass, position, or shape. For example, mass-only guidance produces a narrow peak in the mass marginal while the other morphology metrics remain broad, whereas applying all geometric losses collapses all marginals to narrow peaks at their target values.
@@ 5.7 Geometric Inpainting and Biophysical Simulations
We demonstrate that our geometric guidance framework can controllably edit patient-specific anatomy for simulation experiments. We consider an example involving biventricular pressurization in which we edit a label map to enlarge or shrink the RV. As shown in Figure 11, we define the RV target geometry by doubling or halving the mass measured from the original label map (left column insets). We apply tissue-based inpainting kadry2024probing with geometric guidance to edit the RV (left column) and convert the label map to a tetrahedral mesh (middle column). We simulate biventricular pressurization for the baseline patient and edited variants, showing how RV volume modulates wall displacement (right column). Further details can be found in Section 8.7.
@@ 5.8 Generality over Anatomical Systems and Structures
We aim to show that geometric guidance can control unconditional diffusion models trained on a wide variety of anatomical systems. To this end, we construct and utilize three datasets of 3D multi-class patient-specific label maps: one for the branched ascending aorta, one for the spinal vertebral column, and one for the distal femur comprising the femoral condyles and their articular cartilage, without the patella or tibia. Further details can be found in Section 8.2. In Figure 12, we show typical unconditional samples which do not correspond to the target geometry, as well as samples from geometric guidance, which controls substructure geometry to a high degree of fidelity.
Figure 12: Geometric guidance can control the generation of a wide variety of anatomical systems. We present label maps that were generated from a unconditional and guided latent diffusion models. We train a separate diffusion model for each anatomical dataset. For the aortic dataset, we control the main trunk, for the spinal dataset, we control the sixth, seventh, and eighth thoracic vertebrae, while for the knee dataset, we control the femur.
@@ 6 Limitations
Our method has several limitations, for example, the relative weights of the geometric moments should be obtained through experimental tuning, similar to all guidance frameworks. However, we found that the same set of loss weightings transfer well to entirely different anatomical systems such as the aorta, spinal column, and knee, indicating that only minimal additional tuning is required. Moreover, substructures are currently defined based on label map class, and cannot represent sub-class localized features such as cross sections. This can be addressed by developing localized substructure selection methods, which would enable the control of localized geometric attributes. Lastly, anatomical diffusion models can generate topologically incorrect substructures-such as disconnected aortas or several left atria, making the resulting simulation physics inaccurate. This can be addressed by filtering out topologically incorrect anatomies, at the cost of some wasted computation.
@@ 7 Conclusions
8 Appendix
8.1 Overview
  • In Section 8.2, we provide details on dataset curation and processing.
  • In Section 8.3, we provide implementation details for our autoencoder and diffusion model.
  • In Section 8.4, we provide implementation details for our guidance algorithm.
  • In Section 8.5, we provide implementation details for our conditional generation baselines.
  • In Section 8.6, we provide further experimental details for evaluation and inference.
  • In Section 8.7, we provide implementation details for our biomechanical simulations.
  • In Section 8.8 we provide a dataset scaling analysis for our latent diffusion model.
  • In Section 8.9 we provide qualitative results for a procedurally generated ellipsoid dataset.
  • In Section 8.10 we provide an autoencoder reconstruction error analysis.
  • In Section 8.11 we provide a scale-factor sweep analysis for target modification.
  • In Section 8.12 we provide quantitative results demonstrating disentangled generation for alternative geometric features derived from the second-order moment.
  • In Section 8.13, we present additional morphological distribution plots that examine the effect of guidance weight as well as the choice of control technique.
8.2 Datasets
For our study, we construct four separate datasets of anatomical segmentations to qualitatively demonstrate the flexibility of geometric guidance. These datasets represent 1) whole-heart cardiac segmentations with great vessels, 2) the branched ascending aorta, 3) multi-vertebral spinal column, and 4) the femoral condyle and articular cartilage. We primarily use the cardiac dataset for our experiments.
For the cardiac dataset, we utilize TotalSegmentator v2 wasserthal2022totalsegmentator, with 596 cases manually selected based on segmentation quality assessment. Cardiac structures including the myocardium (Myo), left and right atria (LA & RA), left and right ventricles (LV & RV), aorta (Ao), and pulmonary artery (PA) were segmented using a specialized TotalSegmentator model trained on sub-millimeter resolution data. For the inferior vena cava (IVC), superior vena cava (SVC), and pulmonary veins (PV), we retain the labels from the original dataset. To ensure anatomical validity, we perform topological filtration on all structures except the pulmonary veins, where filtration involves extracting only the largest connected component. The resulting segmentations are standardized by resampling to a uniform voxel resolution of $2\text{mm}$ and subsequently cropped to a fixed range. The crop center is determined from the union of all four chamber segmentations, and the crop size is $128^{3}$ voxels.
For the aorta dataset, we extract labels directly from the original TotalSegmentator v2 wasserthal2022totalsegmentator segmentations, without applying a specialized model, resulting in 450 3D segmentations manually selected based on segmentation quality assessment. The labels include the main aortic trunk and the ascending branches, which comprise the brachiocephalic trunk (BCT), left common carotid artery (LCCA), right common carotid artery (RCCA), left subclavian artery (LSCA), and right subclavian artery (RSCA), for a total of 7 channels per segmentation. All segmentations are resampled to an isotropic voxel size of $2\,\mathrm{mm}$ and cropped to a spatial size of $128^{3}$ using a crop center determined from the center of all combined tissues.
For the spinal dataset, we utilize the CTSpine1K dataset deng2021ctspine1k and extract all vertebral body segmentations, resulting in 784 3D segmentations. The segmentations include 7 cervical vertebrae (C1–C7), 12 thoracic vertebrae (T1–T12), and 5 lumbar vertebrae (L1–L5), for a total of 25 channels per segmentation. To ensure spatial consistency and anatomical completeness, all segmentations are first resampled to an isotropic voxel spacing of $1\,\mathrm{mm}$. The center of the crop box is determined from the union (voxelwise sum) of all vertebral structures in each scan, and a fixed crop of $128^{3}$ voxels is applied for each patient.
For the knee dataset, we utilize the ShapeMedKnee dataset gatti2024shapemed and extract 2000 3D segmentations of the left knee. The segmentations include the femur (Fe) and articular cartilage (Ca), resulting in 3 channels per segmentation. To ensure spatial consistency and anatomical completeness, all segmentations are first resampled to an isotropic voxel spacing of $1\,\mathrm{mm}$. A fixed crop size of $128^{3}$ voxels is applied for each patient.
8.3 Latent Diffusion Model Implementation
For this study, we adapt the VAE and LDM architectures specified by kadry2024probing. The VAE input and output channel counts are set to 11, corresponding to 10 distinct cardiac labels along with an additional channel for the background. The number of input channels for the LDM is set to 3 for unconditional sampling. The hyperparameters and training configuration for the VAE and LDM are listed in Table 2 and Table 3 respectively.
Table 2: Autoencoder hyperparameters
Hyperparameter Value
lr $1\times 10^{-5}$
Epochs 40
Batch Size 1
Num. Channels [16,32,64]
Num. Res. Blocks 2
Downscaling Factor 4
Recon. Loss Weight 1
KL Weight $1\times 10^{-6}$
Table 3: Diffusion model hyperparameters
Hyperparameter Value
Training
lr $2.5\times 10^{-5}$
Epochs 50
Batch Size 1
Num. Channels [64, 128, 196]
Num. Res. Blocks 2
Num. Attn. Heads 1
Attn. Res. 8, 4, 2
$\sigma_{\text{data}}$ 1
$p(\sigma)$ mean 1
$p(\sigma)$ std 1.2
Sampling
$\sigma_{min}$ $1\times 10^{-2}$
$\sigma_{max}$ 80
$\rho$ 3
8.4 Geometric Guidance Implementation
8.4.1 Geometric Moment Computation
To ensure that the extracted components yield interpretable moments, we require the voxel grid values to be softly binarized, with one tissue channel approaching 1 while the others are close to 0. To achieve this, we apply a softmax function with a temperature of $1$. During the computation of geometric moments, we observed that segmentations that are empty or nearly empty, particularly those with small components, lead to unstable gradients that significantly degrade the quality of generation. This instability arises because the centroid and covariance loss calculations utilize mass in the denominator. To mitigate this issue, we introduce a small amount of noise to the mass term whenever it appears in the denominator, thereby stabilizing the overall process. After computing all moments, we normalize the mass term by the total number of voxels $N=HWD$ such that the term represents volume fraction. Unless stated otherwise, we use 50 denoising steps.
8.4.2 Guidance Weight Tuning
We determine the weight factors $\lambda=[\lambda_{0},\lambda_{1},\lambda_{2}]$ for our geometric loss through tuning each loss in isolation. We tune for conditional fidelity while retaining reasonable generation quality metrics. The final weight values can be seen in Table 4.
Table 4: Geometric moment losses and their corresponding weight factors.
Guidance Loss Weight Factor $\lambda$
$\mathcal{L}_{\text{size}}$ $10^{7}$
$\mathcal{L}_{\text{pos}}$ $10^{5}$
$\mathcal{L}_{\text{shape}}$ $10^{4}$
8.5 Baseline Methods Implementation
  • Explicit Conditioning: To ensure that the elements of $\mathcal{G}_{\text{exp}}$ are roughly between 0 and 1, we min-max normalize the masses $\mathcal{M}$, centroids $\mathcal{C}$, and normalized covariances $\mathcal{S}_{n}$ with values calculated from the real dataset (Table 5). The LDM input channel count is increased to accommodate the concatenated input. This method does not readily permit the use of dropout to train a diffusion model in an unconditional manner because the null condition is defined as zero—equivalent to the minimum moment values. We include explicit conditioning results for guidance weights smaller than 0 in Figure 4 for completeness.
  • Cross-Attention Conditioning: Our initial tokens consist of 13-dimensional vectors representing the concatenation of mass $\mathcal{M}$, centroids $\mathcal{C}$, and normalized covariances $\mathcal{S}_{n}$. The tokens are then min-max normalized similar to explicit conditioning and embedded into a 256 dimensional vector for cross-attention. To embed the component index, we use a linear embedding layer. To embed the geometric moments, we use an MLP with three linear layers and apply a ReLU operation after the first and second layers. Both embeddings are added together and used to condition the U-Net with cross-attention, where we use 8 attention heads. To enable unconditional generation, we randomly dropout each channel of $\mathcal{G}_{\text{cross}}$ with a probability of 0.1.
  • Implicit Conditioning: To compute the ellipsoidal distance map, we use the centroids $\mathcal{C}$ and non-normalized covariances $\mathcal{S}$ for each component to compute the Mahalanobis distance de2000mahalanobis for each voxel position. We then apply a shifted sigmoid transform—with a slope of -0.5 and a bias of 1 to constrain the outputs between 0 and 1, and subsequently concatenate the resulting grid to the latents. To enable unconditional generation, we randomly dropout each channel of $\mathcal{G}_{\text{imp}}$ with a probability of 0.1. One limitation of this approach is that the target mass can only be targeted indirectly through the non-normalized covariance term, which can be seen in the conditional fidelity plot for size in Figure 4.
Table 5: Normalizing constants for geometric moments during explicit and cross-attention based conditioning.
Geometric Moment Normalizing Minimum Normalizing Maximum
$\mathcal{M}$ $3.19\times 10^{-3}$ $1.3\times 10^{-2}$
$\mathcal{C}$ 0 11
$\mathcal{S}$ $\times 10^{-4} 1×1$$\times 10^{-2}$
8.6 Additional Experimental Details
  • Morphological evaluation metrics: To compute the morphological metrics, the features are normalized by the mean and standard deviation of the real data. To calculate precision and recall, we use 5 neighbors.
  • Pointcloud evaluation metrics To compute the point cloud metrics, we calculate MMD, COV, and NNA for every tissue label using 256 points sampled using farthest point sampling. The metrics are then averaged over the number of components. To compute the pointcloud distances, we approximate Earth Mover’s Distance (EMD) through the Sinkhorn divergence feydy2019interpolating.
  • Disentangled Generation: Disentangled generation is done by zeroing out the inactive loss weights. Exact configuration details are shown in Table 6. We use 50 denoising steps for all generated samples.
  • Compositional Generation: Our compositional generation experiments vary the number of constrained substructures. The exact labels used for each experiment are detailed in Table 7. We use 100 denoising steps for all generated samples.
Table 6: Configuration details for the disentangled generation ablation study. Checkmarks $\checkmark$ indicate the associated weight factor $\lambda_{i}$ is active while $\times$ indicates the weighting factor is zeroed out.
Guidance Loss $\lambda_{0}$ $\lambda_{1}$ $\lambda_{2}$
Uncond. $\times$ $\times$ $\times$
$\mathcal{L}_{\text{size}}$ $\checkmark$ $\times$ $\times$
$\mathcal{L}_{\text{pos}}$ $\times$ $\checkmark$ $\times$
$\mathcal{L}_{\text{shape}}$ $\times$ $\times$ $\checkmark$
$\mathcal{L}_{\text{geom}}$ $\checkmark$ $\checkmark$ $\checkmark$
Table 7: Configuration details for the compositional generation study.
Substructures Labels
0 None
1 RV
2 RV, RA
3 RV, RA, PA
6 RV, RA, PA, LV, LA, Ao
8.7 Biomechanical Simulation Details
  • Biventricular Cropping: As only myocardial tissue is available for the left ventricle, we approximate an RV myocardial wall by dilating the RV cavity mask to a constant thickness of 4 mm (2 voxels) corresponding to the clinical literature ho2006anatomy. To crop the left and right ventricles at the base of the heart, we define a vector from the LV centroid to the LA centroid, and crop the ventricles by adjusting the position threshold along the defined direction.
  • Tetrahedral Meshing and Processing: The segmentation is then converted into a surface mesh using marching cubes, with a voxel size of 2 mm. Tetrahedral mesh generation is performed using the open-source software Gmsh and MeshLab. The three anatomical models, large RV, baseline patient, and small RV (see Figure 11), are discretized into 39,780, 42,768, and 47,347 linear tetrahedral elements, respectively, with an average edge length of 2 mm.
  • Pressurization Simulation: An in-house finite element method (FEM) solver, implemented in Fortran with MPI, is used for the simulations. The solver is based on the variational multiscale method, providing stabilized FEM formulations GORAYA2024100105; kang2022variational. Simulation results are visualized using the open-source package ParaView.
    The myocardium is modeled as a standard neo-Hookean material with a Young’s modulus of 25 kPa and a Poisson’s ratio of 0.4. Physiological pressure loads of 12 mmHg and 6 mmHg were applied to the LV and RV endocardium, respectively, corresponding to normal diastolic blood pressure. To constrain rigid body motion, zero-displacement Dirichlet boundary conditions are imposed at the base of the heart, while a stress-free Neumann boundary condition is applied on the pericardium.
    The nonlinear finite-deformation elasticity problem is then solved using the Newton–Raphson (NR) method. A direct solver (MUMPS) is employed to solve the discretized algebraic system at each NR iteration, with a convergence tolerance set to $10^{-20}$ for the initial residual. Simulations are carried out on a cluster using 128 processors.
8.8 Dataset Scaling Analysis
We aim to understand the effect of dataset size on both generation quality and conditional fidelity under guidance. To this end, we train four additional autoencoder–diffusion model pairs at different split sizes, using 20%, 40%, 60%, and 80% of the original training set, while keeping the validation set fixed across all models. As shown in Figure 13, generation-quality metrics improve as the training dataset size increases up to 40%. In contrast, conditional fidelity under geometric guidance remains approximately invariant across dataset sizes.
Figure 13: Conditional fidelity is invariant to training set size, while generation quality metrics benefit from training set size up to a limit. Line plots show conditional fidelity and generation quality for latent diffusion models trained on different-sized datasets. In this plot, the right ventricle is constrained.
8.9 Parametric Ellipsoid Dataset Analysis
To further characterize our geometric guidance procedure in isolation from complex anatomical variation, we construct a toy dataset of 3D two-channel ellipsoidal label maps with varying sizes, shapes, and positions. To generate each voxel map, we sample the ellipsoidal radii uniformly from 0 to $0.5$, where $1$ corresponds to the full length of the voxel map. We additionally sample Euler angles uniformly from 0 to $2\pi$, and choose the centroid to lie anywhere within the voxel map such that the ellipsoid is not cropped by the voxel boundaries. We generate $800$ training and $200$ validation label maps.
We then train a latent diffusion model with double the number of base channels to accommodate the large geometric variation in the dataset. We apply our geometric guidance method with centroid and covariance loss weights multiplied by a factor of 10. As shown in Figure 14, our geometric guidance framework can enforce precise geometric constraints on parametric ellipsoid geometries in a disentangled manner.
Figure 14: Geometric guidance can control unconditional diffusion models of procedurally generated ellipsoids. We generate ellipsoidal label maps with varying loss-function combinations to achieve disentangled control.
8.10 Autoencoder Reconstruction Fidelity Analysis
We aim to determine whether the conditional fidelity metrics and topological quality are lower-bounded by the VAE reconstruction error. We first auto-encode 24 seed label maps and measure conditional fidelity for size, position, and shape, as well as the Betti error for each anatomical structure. We then sample 24 label maps over 50 diffusion steps with and without right-ventricular geometric guidance. As summarized in Table 8, geometric guidance substantially improves conditional fidelity relative to unconditional sampling, while the resulting errors for position and shape remain above the VAE reconstruction error.
In terms of topology, we quantify quality using the Betti error, defined as the number of extra connected components relative to the expected topology (e.g., if the aorta is expected to be a single connected component but two are measured, the Betti error is 1). We observe that the VAE introduces only a small number of topological defects, whereas the unconditional diffusion model produces more frequent errors, especially for the aorta (Ao) and pulmonary artery (PA) labels. Finally, we find that geometric guidance can further increase the Betti error, particularly for the PA and inferior vena cava (IVC) labels.
Table 8: Conditional fidelity and Betti error rates for reconstructed or synthetic label maps. Betti error is computed as the mean number of connected components minus 1. Values for size, position, and shape fidelity were multiplied by 1e5, 1e4, 1e4 respectively.
Cond. Metrics Connected Component Betti Error
Method Size Pos. Shape Ao PA IVC SVC LA RA LV RV Myo
VAE Recon. $\mathbf{4.74}$ $\mathbf{2.54}$ $\mathbf{21.53}$ $\mathbf{0.125}$ $\mathbf{0.0}$ $\mathbf{0.125}$ $0.042$ $\mathbf{0.042}$ $\mathbf{0.083}$ $\mathbf{0.0}$ $0.083$ $\mathbf{0.0}$
Unconditional $106.03$ $168.75$ $512.42$ $1.208$ $0.583$ $\mathbf{0.125}$ $\mathbf{0.0}$ $0.083$ $\mathbf{0.083}$ $\mathbf{0.0}$ $\mathbf{0.0}$ $\mathbf{0.0}$
Guided $9.34$ $25.61$ $81.77$ $1.458$ $1.042$ $0.417$ $0.042$ $0.083$ $0.333$ $0.042$ $0.083$ $0.042$
8.11 Editing Scale Factor Analysis
We investigate how far the target right-ventricular mass can be scaled while still producing plausible samples. For each editing factor in $\{0.1,0.5,1.0,2.0,4.0\}$, we take 64 seed label maps, compute the RV mass, multiply it by the editing factor, and use the scaled mass as the conditioning target for mass-only geometric guidance. As summarized in Table 9, decreasing the target mass (factors <1<1) yields samples whose size error and distributional metrics remain close to the unedited case ($\text{factor}=1$): size error increases moderately, and FMD and 1-NNA remain within the same order of magnitude as the baseline. In contrast, increasing the target mass beyond a factor of 2 leads to clear degradation: at a factor of 4, both the size error and FMD increase by more than one order of magnitude, and 1-NNA worsens, indicating that strong mass upscaling produces distorted label maps.
Table 9: We generate label maps with mass-only geometric guidance applied to the right ventricle and artificially changing the target mass derived from the seed label map.
Cond. Metrics Morph. Metrics Pointcloud Metrics
Editing Factor Size FMD ($\downarrow$) Pr. ($\uparrow$) Re. ($\uparrow$) MMD ($\downarrow$) COV ($\uparrow$) 1-NNA
0.1 84.06 126.0 0.14 0.34 13.51 0.295 0.826
0.5 19.95 59.6 0.55 0.50 10.51 0.406 0.713
1.0 12.60 40.3 0.59 0.73 9.26 0.520 0.567
2.0 18.55 167.5 0.16 0.84 11.08 0.430 0.745
4.0 732.10 1690.0 0.00 0.97 26.81 0.273 0.926
8.12 Geometric Guidance with Alternative Moment-features
In our main study, we demonstrated guidance by targeting the normalized second moment to control shape and orientation independently from size. We aim in this section to preliminarily demonstrate that we can achieve fine-grained disentangled control of second-moment derived attributes such as extent, stretch, and orientation. We first decompose the covariance matrix as follows:
$\mathcal{S}=v\mathbf{U}\Lambda^{n}\mathbf{U}^{T},$ (6)
where we define the extent $v\in\mathbb{R}$ as the trace of the eigenvalue matrix $\Lambda\in\mathbb{R}^{3\times 3}$, and normalize $\Lambda$ by $v$ to obtain the anisotropic stretch $\Lambda^{n}=\Lambda/v$. Finally, orientation is represented by the eigenvectors $\mathbf{U}\in\mathbb{R}^{3\times 3}$ derived from the decomposition.
We then define three new geometric losses, which consist of MSE losses for extent and stretch, as well as a dot product loss for orientation.
$\mathcal{L}_{\text{extent}}=\mathcal{L}_{\text{MSE}}(v,\bar{v}),\quad\mathcal{L}_{\text{stretch}}=\mathcal{L}_{\text{MSE}}(\Lambda^{n},\bar{\Lambda}^{n}),\quad\mathcal{L}_{\text{orient}}=\mathcal{L}_{\text{dot}}(\mathbf{U},\bar{\mathbf{U}}).$ (7)
The dot product loss $\mathcal{L}_{\text{dot}}$ is computed as the mean misalignment between corresponding eigenvectors from $\mathbf{U}$ and $\bar{\mathbf{U}}$,
$\mathcal{L}_{\text{dot}}(\mathbf{U},\bar{\mathbf{U}})=\frac{1}{3}\sum_{i=1}^{3}\bigl(1-|\mathbf{u}_{i}^{\top}\bar{\mathbf{u}}_{i}|^{2}\bigr),$ (8)
where $\mathbf{u}_{i}$ and $\bar{\mathbf{u}}_{i}$ denote the $i$-th columns of $\mathbf{U}$ and $\bar{\mathbf{U}}$, respectively, and the absolute value enforces sign-invariance of eigenvector alignment.
With these losses, we conduct a disentangled generation experiment where we sample 32 label maps for each loss ablation setting, with the loss weightings detailed in Table 10. Conditional fidelity for extent and stretch is quantified using the mean absolute error, while conditional fidelity for orientation is quantified using the dot-product loss directly. As shown in Table 11, geometric guidance based on second-order derived features can be applied in a disentangled manner. For example, orientation-only guidance achieves a smaller orientation error while maintaining extent and stretch fidelity comparable to unconditional sampling.
Table 10: Second order moment losses and their corresponding weight factors.
Guidance Loss Weight Factor $\lambda$
$\mathcal{L}_{\text{extent}}$ $10^{5}$
$\mathcal{L}_{\text{stretch}}$ $10^{4}$
$\mathcal{L}_{\text{orient}}$ $10^{2}$
Table 11: We enable disentangled control over geometric features derived from decomposing the second moment into extent, stretch, and orientation. Conditional fidelity metrics for extent and stretch, as well as MMD values were multiplied by 1e3.
Cond. Metrics Morph. Metrics Pointcloud Metrics
Method Extent Stretch Orient. FMD ($\downarrow$) Pr. ($\uparrow$) Re. ($\uparrow$) MMD ($\downarrow$) COV ($\uparrow$) 1-NNA
None 2.14 42.30 0.21 67.63 0.44 0.80 10.22 0.484 0.563
Extent Only 1.42 41.85 0.20 61.99 0.50 0.78 10.12 0.459 0.566
Stretch Only 2.26 3.60 0.23 60.53 0.53 0.80 10.08 0.491 0.559
Orient Only 1.94 41.29 0.0064 60.81 0.59 0.77 10.08 0.525 0.564
8.13 Morphological Analysis
We represent size as the mass of each substructure. Position is represented by the centroid x-coordinate. To characterize shape, we extract the largest eigenvalue and its associated eigenvector from the covariance matrix. Orientation is represented by the polar angle of the principal axis (in spherical coordinates), while elongation is defined as the ratio between the largest eigenvalue and the second-largest eigenvalue.
Additional morphological plots are presented below. In Figure 16, we show that geometric guidance better aligns the distribution of geometric features when comparing real and synthetic anatomies. Figure 15 shows that all geometric-control methods can recapitulate the morphological distribution exhibited by the real data.
Figure 15: Geometric guidance can help recapitulate morphological distributions. Pair plot shows kernel density estimate plots (diagonals) and pairwise scatterplots (off-diagonals) for various morphological metrics. We plot metrics for anatomies generated through conditional baselines and geometric guidance (ours). In this plot, the myocardium is being constrained.
Figure 16: Geometric guidance improves morphological distribution similarity between real and synthetic anatomy. Pair plot shows morphological relationships for mass (top left panel), centroid (bottom left panel), normalized axis lengths (top right panel), and orientation (bottom right panel), where the myocardium labels are being constrained. Diagonal plots show kernel density estimates (LV vs RV), off-diagonal plots show pairwise scatterplots.

### v3

@@ 5.3 Evaluation Metrics
We evaluate pairwise conditional fidelity for size, shape, and position by taking the $L_{1}$-norm between the target and sample moments, averaging over all relevant substructures. We measure morphological quality metrics by comparing the distribution of real and synthetic anatomy in morphological feature space (Kadry et al., 2024). To embed each label map, we consider all 10 tissues as substructures and concatenate, over all substructures, the masses, centroids, and eigenvalues of the normalized covariance matrices. We specifically use morphological variants of improved precision and recall, as well as the Fréchet distance (FD) (Kynkäänniemi et al., 2019; Kadry et al., 2024). Lastly, we leverage pointcloud-based metrics to assess 3D shape (Yang et al., 2019), such as minimum matching distance (MMD), coverage (COV), and 1-nearest neighbor accuracy (1-NNA). Distances between pointclouds are computed with Earth Mover’s Distance (EMD). Further details can be found in Section 9.6.
@@ 5.4 Evaluating Anatomical Generation Quality
We first aim to compare and evaluate geometric control methods on both conditional fidelity and synthetic anatomy quality. We sample target moments for a single substructure (myocardium) from the validation set and generate 200 anatomical segmentations per method. We sweep over guidance weights $w\in[0,2]$. In Figure 4, we show that geometric guidance enhances conditional fidelity, especially at higher guidance weights. We observe that our method maintains generation quality, retaining similar levels of morphological and pointcloud evaluation metrics with increasing guidance. Figure 3 shows example label maps generated through varying guidance values for all methods; only our method and implicit conditioning align the target and sample ellipsoids under guidance. Further information on which features were plotted can be found in Section 9.13.
@@ 5.5 Evaluating Geometric Disentanglement
We next show that our guidance framework uniquely enables disentangled control of geometric attributes. We use 100 target moments for myocardial labels from the validation set using no losses (Uncond.), a combination of all losses ($\mathcal{L}_{\text{geom}}$), or each individual moment loss ($\mathcal{L}_{\text{size}},\mathcal{L}_{\text{pos}}$, and $\mathcal{L}_{\text{shape}}$). Figure 9 shows that each individual loss improves its corresponding conditional-fidelity metric while leaving the others approximately unchanged. The main exception is the interaction between shape and mass, where adding a guidance weight for the shape loss enhances mass fidelity. This phenomenon is likely due to correlations between size and shape in the dataset. Qualitative results can be seen in Figures 5 and 6, where the right ventricle is constrained independently by mass, position, or shape. For example, mass-only guidance produces a narrow peak in the mass marginal while the other morphology metrics remain broad, whereas applying all geometric losses collapses all marginals to narrow peaks at their target values.
@@ 5.7 Geometric Inpainting and Biophysical Simulations
We demonstrate that our geometric guidance framework can controllably edit patient-specific anatomy for simulation experiments. We consider an example involving biventricular pressurization in which we edit a label map to enlarge or shrink the RV. As shown in Figure 11, we define the RV target geometry by doubling or halving the mass measured from the original label map (left column insets). We apply tissue-based inpainting (Kadry et al., 2024) with geometric guidance to edit the RV (left column) and convert the label map to a tetrahedral mesh (middle column). We simulate biventricular pressurization for the baseline patient and edited variants, showing how RV volume modulates wall displacement (right column). Further details can be found in Section 9.7.
@@ 5.8 Generality over Anatomical Systems and Structures
We aim to show that geometric guidance applies to a wide variety of anatomical diffusion models. We construct three datasets of 3D multi-class patient-specific label maps corresponding to the branched ascending aorta, spinal vertebral column, and distal femur. Further details can be found in Section 9.2. In Figure 12, we show typical unconditional samples, as well as samples from geometric guidance, which constrains substructure geometry to a high degree of fidelity.
Figure 12: Geometric guidance can control the generation of a wide variety of anatomical systems. We present label maps that were generated from guided and unguided latent diffusion models. For the aortic dataset, we control the main trunk, for the spinal dataset, we control the sixth, seventh, and eighth thoracic vertebrae, while for the knee dataset, we control the femur.
@@ 6 Limitations
Our method has several limitations. First, the relative weights of the geometric moments should be obtained through experimental tuning, similar to all guidance frameworks. However, we found that the same set of loss weightings transfer well to entirely different anatomical systems such as the aorta, spinal column, and knee, indicating that only minimal additional tuning is required. Second, substructures are currently defined based on label map class, and cannot represent localized features such as cross sections. Lastly, anatomical diffusion models can generate topologically incorrect substructures, such as disconnected aortas or several left atria, making the resulting simulation physics inaccurate. This can be addressed by filtering out topologically incorrect anatomies, at the cost of some wasted computation.
@@ 7 Conclusions
8 Acknowledgments
FRN gratefully acknowledges support from the American Heart Association Career Award (25CDA1452999).